OpenStudy (anonymous):

what can i use in a sentence to describe the change that is happening in a linear function

OpenStudy (isaiah.feynman):

Simple. The change is constant.

OpenStudy (mathmale):

Let's be more precise. In the linear function y = m*x + b, m represents the (constant) slope. Thus, any change in x, multiplied by the slope, m, results in the corresponding change in y.

OpenStudy (mathmale):

Example 1: if y = 3x - 2, and x increases by 1, then y increases by 3(1), or 3. Example 2: if y = (-1/2) + 7, and x increases by 1, then y decreases by (-1/2)(1) = -1/2.